1.  A flask contains 2 grams of helium at a pressure of 12 atmospheres and a temperature of 57oC. (a) What is the volume of the flask?. The system is checked at a later time when the temperature is 17oC and the pressure is found to be only 7 atmospheres. How many grams of helium have leaked out of the flask?

 



1
Expert's answer
2021-02-06T10:04:19-0500

a)


"pV=\frac{m}{M}RT\to V=\frac{m}{pM}RT="


"=\frac{2\cdot10^{-3}}{12\cdot 101325\cdot 4\cdot10^{-3}}\cdot8.31\cdot 330=0.00112\ (m^3)" .Answer



b)


"pV=\frac{m}{M}RT\to m=\frac{pVM}{RT}="


"=\frac{7\cdot 101325 \cdot 0.00112\cdot 0.004}{8.31\cdot 290}=0.00132\ (kg)"


"\Delta m=0.002-0.00132=0.00068\ (kg)=0.68\ (g)" .Answer